OpenStudy (anonymous):

How do I factor x^2-3x+5=0 ? I have looked at it and I keep trying and I get nowhere! Please help.

OpenStudy (abb0t):

Use the quadratic formula: \[\frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] where \[ax^2 +bx + c = 0\]

OpenStudy (anonymous):

Thank you! I was getting no where trying to factor it!